I. The values for the change in entropy and enthalpy was obtained for an unknown chemical reaction.

The reaction started at 10°C and the values for the thermodynamic parameters began to be consistent

at room temperature.


2. The reaction gives the following values at 25°C: ΔH = 2570 J, ΔS = 75 J/K. Determine if this is a

spontaneous reaction.


1
Expert's answer
2021-03-23T07:25:44-0400

The reaction gives the following values at 25°C:

ΔH = 2570 J

ΔS = 75 J/K

ΔG = ΔH – TΔS

T = 273 + 25 = 298 (K)

ΔG = 2570 J – 298 K x 75 J/K = -19780 J < 0 so, this is a spontaneous reaction.